# FitzHugh-Nagumo equation: from global dynamics to the slow-fast system

### v1

@@ 2.3. Poincaré Compactification
In this section, we present a concise overview of the Poincaré compactification following Chapter 5 of [5], a technique used in Section 5 of the present paper. Consider the polynomial vector field (2.2) in $\mathbb{R}^{2}$ where $d_{1}$ and $d_{2}$ are the algebraic degrees of $P$ and $Q$ respectively and $d={\max}\{d_{1},d_{2}\}.$
@@ 2.3.1. Preliminaries for the construction
First identify $\mathbb{R}^{2}$ with the plane $\Pi$ in $\mathbb{R}^{3}$ defined by $(y_{1},y_{2},y_{3})=(x_{1},x_{2},1).$ The sphere $\mathbb{S}^{2}=\{(y_{1},y_{2},y_{3})\in\mathbb{R}^{3}:y_{1}^{2}+y_{2}^{2}+y_{3}^{2}=1\}$ is called the Poincaré sphere and is tangent to the plane $\Pi$ at $(0,0,1).$ The sphere may be written as
@@
The projection of $\mathbf{X}$ from $\mathbb{R}^{2}$ to $\mathbb{S}^{2}$ is given by the central projections
$f^{+}:\mathbb{R}^{2}\to\mathbb{S}^{2}\quad\text{and}\quad f^{-}:\mathbb{R}^{2}\to\mathbb{S}^{2}$
where $f^{+}(x_{1},x_{2})$ is the intersection point of the line passing through the origin and $(x_{1},x_{2})\in\mathbb{R}^{2}$, with $H^{+}$. More precisely:
@@
We obtain vector fields in each hemisphere (analytically conjugate to the initial vector field $\mathbf{X}$. For $x=(x_{1},x_{2})\in\mathbb{R}^{2}$, the induced vector field on $H^{+}$ and $H^{-}$ is defined, respectively, by:
$\overline{\mathbf{X}}(y)$ $=Df^{+}(x)\mathbf{X}(x),\quad\text{where}\quad y=f^{+}(x),$
$\overline{\mathbf{X}}(y)$ $=Df^{-}(x)\mathbf{X}(x),\quad\text{where}\quad y=f^{-}(x).$
@@
The vector field $\overline{\mathbf{X}}$ on $\mathbb{S}^{2}\backslash\mathbb{S}^{1}$ is everywhere tangent to $\mathbb{S}^{2}.$
The points at infinity of $\mathbb{R}^{2}$ (each direction is associated with two points) are in bijective correspondence with the points of the equator of $\mathbb{S}^{2}.$ In the next subsection, we are going to extend the induced vector field $\overline{\mathbf{X}}$ from $\mathbb{S}^{2}\backslash\mathbb{S}^{1}$ to $\mathbb{S}^{2}$.
@@
The extended vector field on $\mathbb{S}^{2}$ is called the Poincaré compactification of the vector field $\mathbf{X}$ on $\mathbb{R}^{2},$ and it is denoted by $p(\mathbf{X}).$
@@ 2.3.2. About the construction
We use smooth charts to make calculations. For $y=(y_{1},y_{2},y_{3})\in\mathbb{S}^{2},$ we use the six local charts given by
$U_{k}=\{y\in\mathbb{S}^{2}:y_{k}>0\},\quad V_{k}=\{y\in\mathbb{S}^{2}:y_{k}<0\},\quad k=1,2,3.$
The corresponding local maps $\Phi_{k}:U_{k}\to\mathbb{R}^{2}$ and $\Psi_{k}:V_{k}\to\mathbb{R}^{2}$ are defined as
$\Phi_{k}(y)=-\Psi_{k}(y)=\left(\frac{y_{m}}{y_{k}},\frac{y_{n}}{y_{k}}\right),\quad\text{for}\quad m<n\quad\text{and}\quad m,n\neq k.$
We denote by $z=(u,v)$ the value of $\Phi_{k}(y)$ or $\Psi_{k}(y)$ for any $k\in\{1,2,3\}$, such that $(u,v)$ will play different roles depending on the local chart we are considering. Observe that points lying in $\mathbb{S}^{1}$, in any chart, have $v=0$.
We perform a detailed calculation of the expression of $p(\mathbf{X})$ in the local chart $U_{1}$. From (2.2), we have $\mathbf{X}(x)=(P(x_{1},x_{2}),Q(x_{1},x_{2}))$. Then $\overline{\mathbf{X}}(y)=Df^{+}(x)\mathbf{X}(x)$ with $y=f^{+}(x)$ and, using the Chain rule, one gets:
$D\Phi_{1}(y)\overline{\mathbf{X}}(y)=D\Phi_{1}(y)\circ Df^{+}(x)X(x)=D(\Phi_{1}\circ f^{+})(x)\mathbf{X}(x).$
Let $\overline{\mathbf{X}}|_{U_{1}}$ denote the system defined as $D\Phi_{1}(y)\overline{\mathbf{X}}(y)$. Then since
$(\Phi_{1}\circ f^{+})(x)=\left(\frac{x_{2}}{x_{1}},\frac{1}{x_{1}}\right)=(u,v),$
we have
$\overline{\mathbf{X}}|_{U_{1}}=v^{2}\left(-\frac{u}{v}P\left(\frac{1}{v},\frac{u}{v}\right)+\frac{1}{v}Q\left(\frac{1}{v},\frac{u}{v}\right),-P\left(\frac{1}{v},\frac{u}{v}\right)\right).$
On the other hand, one knows that $\rho(y)=y_{3}^{d-1}=v^{d-1}m(z)$ where
$m(z)=(1+u^{2}+v^{2})^{\frac{1-d}{2}}.$
Then we get:
$\rho\left(\overline{\mathbf{X}}|_{U_{1}}\right)(z)=v^{d+1}m(z)\left(-\frac{u}{v}P\left(\frac{1}{v},\frac{u}{v}\right)+\frac{1}{v}Q\left(\frac{1}{v},\frac{u}{v}\right),-P\left(\frac{1}{v},\frac{u}{v}\right)\right).$
Remark 2.6.
In order to prove that the extension of $\rho\overline{\mathbf{X}}$ to $p(\mathbf{X})$ is defined on the whole of $\mathbb{S}^{2}$ we notice that while $\overline{\mathbf{X}}|_{U_{1}}$ is not well defined when $v=0$, and $p(\mathbf{X})|_{U_{1}}=\rho\overline{\mathbf{X}}|_{U_{1}}$ is well defined along $v=0$, since the multiplying factor $v^{d+1}$ cancels any factor of $v$ which may appear in the denominator. The same line of arguments may be applied to the other local charts.
To simplify the extended vector field we also make a change in the time variable and remove the factor $m(z)$. We still keep a vector field on $\mathbb{S}^{2}$ which is $C^{\omega}$–equivalent to $\mathbf{X}$ on any of the hemispheres $H^{+}$ and $H^{-}$.
@@ 2.3.3. Explicit expressions
The expression for $p(\mathbf{X})$ in local chart $(U_{1},\Phi_{1})$ is given by:
$\dot{u}=v^{d}\left[-uP\left(\frac{1}{v},\frac{u}{v}\right)+Q\left(\frac{1}{v},\frac{u}{v}\right)\right],\hskip 14.22636pt\dot{v}=-v^{d+1}P\left(\frac{1}{v},\frac{u}{v}\right),$ (2.5)
@@
$\dot{u}=v^{d}\left[P\left(\frac{u}{v},\frac{1}{v}\right)-uQ\left(\frac{u}{v},\frac{1}{v}\right)\right],\hskip 14.22636pt\dot{v}=-v^{d+1}Q\left(\frac{u}{v},\frac{1}{v}\right),$ (2.6)
@@
For $i=1,2,3$, the expressions of the vector field $p(\mathbf{X})$ in the local chart $(V_{i},\Psi_{i})$ are identical to those in the local chart $(U_{i},\Phi_{i}),$ up to a multiplication by $({-1})^{d-1}$.
To study $\mathbf{X}$ in the complete plane $\mathbb{R}^{2}$, including its behaviour near infinity, it suffices to work on $H^{+}\cup\mathbb{S}^{1}$, which we call the Poincaré disc. All calculations can be performed in the three charts $(U_{1},\Phi_{1})$, $(U_{2},\Phi_{2})$, and $(U_{3},\Phi_{3})$ in which case the expressions are given by the formulas (2.5), (2.6) and (2.7). To obtain (2.5) we start with (2.2) and introduce coordinates $(u,v)$ through the equality:
$(x_{1},x_{2})=\left(\frac{1}{v},\frac{u}{v}\right).$
In each local chart, the local representative of $p(\mathbf{X})$ is a polynomial vector field.
@@
We call finite (resp. infinite) singular points of $\mathbf{X}$ the singular points of $p(\mathbf{X})$ which lie in $\mathbb{S}^{2}\backslash\mathbb{S}^{1}$ (resp. $\mathbb{S}^{1}$).
@@ 2.3.4. Useful consequences
    If $y\in\mathbb{S}^{1}$ is an infinite singular point, then $-y$ is also a singular point.
@@
    Since the local behavior near $-y$ is the local behavior near $y$ multiplied by $(-1)^{d-1}$, then:
    1. (a)
      the orientation of the orbits changes when the degree $d$ is even.
    2. (b)
      if $d$ is even and $y\in\mathbb{S}^{1}$ is a stable node of $p(\mathbf{X})$, then $-y$ is an unstable node.
@@
    Infinite singular points appear in pairs of diametrically opposite points so that it is enough to study half of them. Using the degree of the vector field one can determine the other half.
@@
    The integral curves of $\mathbb{S}^{2}$ are symmetric with respect to the origin, such that it is sufficient to represent the flow of $p(\mathbf{X})$ only in the closed northern hemisphere, the so called Poincaré disc.
@@ 2.3.5. Infinite singular points
The theory described here will be useful in Section 5 of the present paper. We aim to study the local phase portrait at infinite singular points. For this we choose an infinite singular point $(u,0)$ and start by looking at the expression of the linear part of the vector field $p(\mathbf{X})$. Denote by $P_{i}$ and $Q_{i}$ the homogeneous polynomials of degree $i\in\mathbb{N}_{0}$ for $i=0,1,...,d$ such that
@@
Then $(u,0)\in\mathbb{S}^{1}\cap(U_{1}\cup V_{1})$ is an infinite singular point of $p(\mathbf{X})$ if and only if
$F(u)\equiv Q_{d}(1,u)-uP_{d}(1,u)=0.$
Similarly $(u,0)\in\mathbb{S}^{1}\cap(U_{2}\cup V_{2})$ is an infinite singular point of $p(\mathbf{X})$ if and only if
$G(u)\equiv P_{d}(u,1)-uQ_{d}(u,1)=0.$
@@
if $(u,0)$ belongs to $U_{1}\cup V_{1}$ or $U_{2}\cup V_{2}$, respectively.
Our discussion is concentrated on isolated singularities in the equator.
Following [5], among the hyperbolic singular points at infinity only nodes and saddles may appear. All the semi-hyperbolic singular points can appear at infinity. If one of these hyperbolic or semi-hyperbolic singularities at infinity is a (topological) saddle, then the straight line defined by $v=0$, representing the equator of $\mathbb{S}^{2}$, is a stable or unstable manifold, or a center manifold. The same property also holds for semi-hyperbolic singularities of saddle-node type. They can have their hyperbolic sectors split in two different ways depending on the Jacobian matrix of the system in the charts $U_{1}$ or $U_{2}$. These matrices may be either
$\left(\begin{array}[]{cc}a&\star\\ \\ 0&0\end{array}\right)\quad\text{or}\quad\left(\begin{array}[]{cc}0&\star\\ \\ 0&a\end{array}\right)$
with $a\neq 0$ and $\star\in\mathbb{R}$. The sense of the orbits can also be the opposite. The nilpotent points, as well as the singularities with zero linear part, have a behavior at infinity that is quite a bit more complicated than the hyperbolic and elementary singular points. Blow-up is needed to study them.
@@ 3. Finite equilibria and bifurcation analysis
$\left\{\begin{array}[]{l}\dot{x}=\displaystyle c\left[y-\left(\frac{x^{3}}{3}-x\right)\right]\\ \\ \dot{y}=\displaystyle-\frac{1}{c}(x-a+by)\end{array}\right.$ (3.1)
where $a,b\in\mathbb{R}$ and $\0c\in\mathbb{R}\backslash\{0\}.$ We can absorb the parameter $z$ of (1.5) into the variable through the change of variable $y\to y+z$ and redefinition of $a\to a-bz.$ The model (3.1) is invariant under the transformation $(t,c)\to(-t,-c)$ so we can restrict to the case $c>0$, which means that the stability of the equilibria for $c>0$ is the reverse of that of $c<0$.
Let us denote by $f_{(a,b,c)}:\mathbb{R}^{2}\to\mathbb{R}^{2}$ the vector field associated with (3.1). For $a,b\in\mathbb{R}$ and $\0c\in\mathbb{R}\backslash\{0\}$, the Jacobian matrix of $f_{(a,b,c)}$ at a general point $(x,y)\in\mathbb{R}^{2}$ is given by:
@@ 3.1. Case A ($a=0$)
For all $a=0$, $b\in\mathbb{R}$ and $\0c\in\mathbb{R}\backslash\{0\},$ we know that $E_{1}=(0,0)$ is an equilibrium of (3.1). If $b\in(-\infty,0)\cup[1,+\infty),$ then system (3.1) has two extra equilibria given explicitly by
@@
It is easy to see that $E_{1}$ is a hyperbolic saddle for $b>1$ (since $\mbox{Det}(E_{1})<0$). Using (3.5), the jacobian matrix of the vector field $f_{(0,1,\pm 1)}$ at $E_{1}\equiv E_{2}\equiv E_{3}$ is given by:
(±$\left(\begin{array}[]{cc}\pm 1&\pm 1\\ \\ \mp 1&\mp 1\end{array}\right).$ (3.10)
It is easy to verify that the matrix (3.10) is non-hyperbolic and has a double zero eigenvalue. Our main result relies on the existence of a Double-zero bifurcation (DZ) with $\mathbb{Z}_{2}(\kappa)$–symmetry of $f_{(0,1,\pm 1)}$ at the singularity $E_{1}$.
@@
so that the equations in Jordan form are given by
@@
which contains part of the truncated form of degree 3 of the versal deformation of a Double-zero (DZ) bifurcation with $\mathbb{Z}_{2}$–symmetry [27, pp. 400, Case $c=-1$]. The next result deals with the description of all bifurcation curves passing through the bifurcation point $(b,c)=(1,\pm 1)$ and that characterises a DZ bifurcation of $f_{(0,b,c)}$ with $\mathbb{Z}_{2}(\kappa)$–symmetry. We focus the proof on the analysis around the bifurcation parameter $(b,c)=\left(1,1\right)$; the analysis close to $(b,c)=\left(1,-1\right)$ has a similar treatment (reversing time).
@@
With respect to the vector field $f_{(0,b,c)},$ the equilibrium $E_{1}\equiv E_{2}\equiv E_{3}$ undergoes a DZ bifurcation with $\mathbb{Z}_{2}(\kappa)$–symmetry at $(b,c)=(1,1).$ The global representations of the transition/bifurcation curves in the space of parameters $(b,c)\in\mathbb{R}\times\mathbb{R}^{+}$ are as follows (schematic lines for $a=0$ have been plotted in Figures 1 and 3):
@@
    The equilibrium $E_{1}$ undergoes a subcritical (with respect to $b$) Hopf bifurcation along part of the parabola:
@@
    The equilibria $E_{2}$ and $E_{3}$ undergo a supercritical (with respect to $b$) Hopf bifurcation along the algebraic curves:
@@
Figure 1. Illustration of Theorem 3.4. Left: The bifurcation curves associated to (3.1) for Case A. Linear stability of $E_{1}$ for $f_{(0,b,c)}$. Right: Linear stability of $E_{2}$ and $E_{3}$ for $f_{(0,b,c)}$. The notation $T_{P}$, $T_{H}^{1}$, $T_{F}^{1}$, $T_{P}$ $T_{H}^{2,3}$, $T_{F}^{2,3}$ follows from Theorem 3.4 and sd, uf, un, sf, and sn refer to saddle, unstable focus, unstable node, stable focus and stable node. The curve $b=0$ corresponds to a subcritical pitchfork bifurcation from where the equilibria $E_{2}$ and $E_{3}$ (3.6) collapse into $E_{1}$.
@@
Theorem 3.4 generalizes the analysis of [8] and [21] by locating a codimension 2 bifurcation with $\mathbb{Z}_{2}(\kappa)$–symmetry at $(b,c)=(1,1)$ for the vector field $f_{(0,b,c)}.$ For the sake of completeness, we plot a computer-assisted global phase portraits on Table 1 of Appendix. Figure 8 sketches the main features of the phase portraits in the different regions – within each region of the figure, the dynamics is $C^{1}$–conjugated to the respective figure.
@@
    We prove the existence of a Pitchfork bifurcation for $b=1$ using the normal form truncated at order 3 [27, Section 20.1E]. Let $b=1+\mu$, where $\mu\in\mathbb{R}$ is an additional bifurcation parameter. For $\mu=0$ the Jacobian matrix (3.5) associated with $E_{1}$ admits eigenvalues: 00 and $\frac{c^{2}-1}{c}$, $c\neq 0,1$. Define the diffeomorphism $T:\mathbb{R}^{2}\to\mathbb{R}^{2}$ as
    $\mathbf{x}\equiv(x,y)\mapsto\left(\frac{c^{2}y-x}{c^{2}-1},\frac{x-y}{c^{2}-1}\right).$
[…]
@@
    We aim to find a center manifold in the form of $h(x,\mu):={\alpha_{0}}^{2}\mu+\alpha_{1}x^{2}+\alpha_{2}x\mu.$ Since the center manifold must satisfy the equation ([27, Equation 3.2.7])
@@
    meaning that there is a Pitchfork bifurcation at $\mu=0$. If $c>1$, then the bifurcation is supercritical with respect to $\mu$ and also with respect to $b$ in the original equation, as depicted in Figure 2. The analysis for the case $c>1$ is different from the case $10<c<$ 1, although the conclusion is the same.
@@
    Let $b=c^{2}$ and $|c|<1$. The Jacobian matrix of $f_{(0,b,c)}$ at $E_{1}$ has a pair of purely imaginary eigenvalues $i\sqrt{1-c^{2}}$ with $-1<c<1.$ Let $b=c^{2}+\mu$ where $\mu\in\mathbb{R}$ is a bifurcation parameter. The eigenvalues associated with the Jacobian matrix at $E_{1}$ for $\mu\neq 0$
@@
    $\begin{pmatrix}u\\ v\end{pmatrix}=\begin{pmatrix}\dfrac{1}{2}&\dfrac{1}{2}\dfrac{c}{\sqrt{1-c^{2}}}\\ 0&-\dfrac{1}{2}\dfrac{1}{c\sqrt{1-c^{2}}}\end{pmatrix}\begin{pmatrix}x\\ y\end{pmatrix},$
@@
    By calculating the first Lyapunov coefficient ([26, Equations (2) and (3)]), we have:
    $l:=\dfrac{1}{16\omega}\left(R_{1}+\omega R_{2}\right),$

### v3

@@ 2.3. Poincaré compactification
Here, following Chapter 5 of [5], we present a concise overview of the Poincaré compactification to study how trajectories of a planar differential system behave “near infinity”. This technique will be used in Section 5 of the present paper.
To construct the phase portrait, one would ideally need to consider the complete real plane $\mathbb{R}^{2},$ which is often impractical. However, if the vector field is defined by polynomial functions, Poincaré compactification can be applied. This method allows the phase portrait to be represented within a finite region and also provides control over orbits that approach or come from infinity. Consider the polynomial vector field $f=(P,Q)$, where $d_{1}$ and $d_{2}$ are the algebraic degrees of $P$ and $Q$ respectively and $d={\max}\{d_{1},d_{2}\}.$
@@ 2.3.1. Preliminaries for the construction
First identify $\mathbb{R}^{2}$ with the plane $\Pi$ in $\mathbb{R}^{3}$ defined as $(y_{1},y_{2},y_{3})=(x_{1},x_{2},1).$ The sphere
$\mathbb{S}^{2}=\left\{(y_{1},y_{2},y_{3})\in\mathbb{R}^{3}:\sum_{i=1}^{3}{y_{i}}^{2}=1\right\}$
is called the Poincaré sphere and is tangent to $\Pi$ at $(0,0,1).$ The sphere may be written as
@@
The projection of $\mathbf{X}:\mathbb{R}^{2}\to\mathbb{S}^{2}$ is defined with the central projections
$f^{\pm}:\mathbb{R}^{2}\to\mathbb{S}^{2}$
where $f^{\pm}(x_{1},x_{2})$ is the intersection point of the line that passes through $(0,0)$ and $(x_{1},x_{2})\in\mathbb{R}^{2}$, with $H^{\pm}$. More precisely:
@@
For $x=(x_{1},x_{2})\in\mathbb{R}^{2}$, the induced vector field on $H^{\pm}$ is defined, respectively, by:
$\overline{\mathbf{X}}(y)$ $=Df^{\pm}(x)\mathbf{X}(x),\quad\text{for}\quad y=f^{\pm}(x).$
@@
The vector field $\overline{\mathbf{X}}$ is tangent to the sphere $\mathbb{S}^{2}$ at every point $\mathbb{S}^{2}\backslash\mathbb{S}^{1}$.
The infinite points of $\mathbb{R}^{2}$ (each direction is associated with two points) are in bijective associated to the points of the equator of $\mathbb{S}^{2}.$ In the next subsection, we are going to generalize the induced vector field $\overline{\mathbf{X}}:\mathbb{S}^{2}\backslash\mathbb{S}^{1}\rightarrow\mathbb{S}^{2}.$
@@
The extension of $\mathbf{X}$ to $\mathbb{S}^{2}$ is referred to as its Poincaré compactification, denoted by $p(\mathbf{X}).$
@@ 2.3.2. About the construction
We use smooth charts to make calculations. For $y=(y_{1},y_{2},y_{3})\in\mathbb{S}^{2},$ we utilize the six local charts:
$U_{l}=\{y\in\mathbb{S}^{2}:0<y_{l}\},\quad V_{l}=\{y\in\mathbb{S}^{2}:0>y_{l}\},\quad l=1,2,3.$
The associated local maps $\Phi_{l}:U_{l}\to\mathbb{R}^{2}$ and $\Psi_{l}:V_{l}\to\mathbb{R}^{2}$ are described as
$\Phi_{l}(y)=-\Psi_{l}(y)=\left(\frac{y_{m}}{y_{l}},\frac{y_{n}}{y_{l}}\right),\quad\text{for}\quad m<n\quad\text{where}\quad m,n\neq l.$
Consider $z=(u,v)$ as the value of $\Phi_{l}(y)$ or $\Psi_{l}(y)$ for any $l\in\{1,2,3\}$, such that $(u,v)$ plays different roles depending on the local chart we are considering. In any chart, points on $\mathbb{S}^{1}$ correspond to those with $v=0$.
@@ 2.3.3. Explicit expressions
In the local chart $(U_{1},\Phi_{1}),$ $p(\mathbf{X})$ is expressed as
$\dot{u}=\left[Q\left(\frac{1}{v},\frac{u}{v}\right)-uP\left(\frac{1}{v},\frac{u}{v}\right)\right]v^{d},\hskip 14.22636pt\dot{v}=-P\left(\frac{1}{v},\frac{u}{v}\right)v^{d+1},$ (2.3)
@@
$\dot{u}=\left[P\left(\frac{u}{v},\frac{1}{v}\right)-uQ\left(\frac{u}{v},\frac{1}{v}\right)\right]v^{d},\hskip 14.22636pt\dot{v}=-Q\left(\frac{u}{v},\frac{1}{v}\right)v^{d+1},$ (2.4)
@@
For $i=1,2,3$, the representations for the vector field $p(\mathbf{X})$ in the local chart $(V_{i},\Psi_{i})$ are identical to those in $(U_{i},\Phi_{i}),$ up to a multiplication by $({-1})^{d-1}$.
To investigate $\mathbf{X}$ throughout the complete plane $\mathbb{R}^{2}$, as well as its behaviour “near infinity”, it is enough to examine it on $H^{+}\cup\mathbb{S}^{1}$, a region known as the Poincaré disc. All calculations can be performed in the three charts $(U_{i},\Phi_{i})$ for $i=1,2,3$ whose expressions are given by the formulas (2.3), (2.4) and (2.5).
@@
We refer to singular points of $\mathbf{X}$ as finite when the corresponding points of $p(\mathbf{X})$ are in $\mathbb{S}^{2}\setminus\mathbb{S}^{1},$ and as infinite when they belong to $\mathbb{S}^{1}$.
@@ 2.3.4. Useful consequences
    Whenever $y\in\mathbb{S}^{1}$ is an infinite singular point, so is y-y.
@@
    The local behaviour near y-y is that near $y$ scaled by $(-1)^{d-1}$.
@@
    Because infinite singular points arise in pairs, it is sufficient to examine just half of them. The rest can be inferred via the vector field’s degree.
@@
    The integral curves on $\mathbb{S}^{2}$ are symmetric aroud the origin, so it suffices to depict the flow of $p(\mathbf{X})$ in the closed northern hemisphere, the so called Poincaré disc.
@@ 2.3.5. Infinite singular points
The theory described here will be useful in Section 5 of the present paper. Our goal is to analyse the local phase portrait near infinite singular points. To this end, we select a singular point $(u,0)$ and begin with examining the linear part of the vector field $p(\mathbf{X})$. Denote by $P_{i}$ and $Q_{i}$ the homogeneous polynomials of degree $i\in\mathbb{N}_{0}$ for i=0,1,,di=0,1,...,d such that
@@
Then $(u,0)\in(U_{1}\cup V_{1})\cap\mathbb{S}^{1}$ is an infinite singular point of $p(\mathbf{X})$ if and only if
$F(u)\equiv-uP_{d}(1,u)+Q_{d}(1,u)=0.$
In the same way $(u,0)\in(U_{2}\cup V_{2})\cap\mathbb{S}^{1}$ is an infinite singular point for $p(\mathbf{X})$ if and only if
$G(u)\equiv-uQ_{d}(u,1)+P_{d}(u,1)=0.$
@@
if $(u,0)$ lies in $U_{1}\cup V_{1}$ or $U_{2}\cup V_{2}$, respectively.
Our discussion is concentrated on isolated singularities in the equator. According to [5], only nodes and saddles can arise as hyperbolic singularities at infinity. All semi-hyperbolic singular points may also emerge at infinity. In the case that one of these singular points at infinity is a (topological) saddle, the line $v=0,$ corresponding to the equator of $\mathbb{S}^{2}$, serves as a stable, unstable, or center manifold. This property also applies to semi-hyperbolic points of saddle-node type, whose hyperbolic sectors may be divided into two different forms depending on the Jacobian matrix in charts $U_{1}$ or $U_{2}.$ These matrices may be either
$\left(\begin{array}[]{cc}0&\star\\ \\ 0&a\end{array}\right)\quad\text{or}\quad\left(\begin{array}[]{cc}a&\star\\ \\ 0&0\end{array}\right)$
where $a\neq 0$ and $\star\in\mathbb{R}$. Nilpotent equilibria exhibit behaviour at infinity that is considerably more complex than those of hyperbolic equilibria. Blow-up techniques of Subsection 2.2 are needed to study them.
@@ 3. Finite equilibria and bifurcation analysis
$\left\{\begin{array}[]{l}\dot{x}=\displaystyle c\left[y-\left(\frac{x^{3}}{3}-x\right)\right]\\ \\ \dot{y}=\displaystyle-\frac{1}{c}(x+by-a)\end{array}\right$. (3.1)
[…]
@@ 3.1. Case A ($a=0$)
[…]
@@
It can be easily shown that $E_{1}$ is a hyperbolic saddle for $b>1$ (since $\mbox{Det}(E_{1})<0$). Using (3.5), the jacobian matrix of the vector field $f_{(0,1,\pm 1)}$ at $E_{1}\equiv E_{2}\equiv E_{3}$ is given by:
$\left(\begin{array}[]{cc}\pm 1&\pm 1\\ \\ \mp 1&\mp 1\end{array}\right).$ (3.10)
It is straightforward to see that the matrix (3.10) is non-hyperbolic and has a double zero eigenvalue. Our main result relies on the existence of a Double-zero bifurcation (DZ) with $\mathbb{Z}_{2}(\kappa)$–symmetry of $f_{(0,1,\pm 1)}$ at the singularity $E_{1}$.
@@
so that the equations in Jordan form are represented by
@@
which contains part of the truncated form of degree 3 of the versal deformation of a Double-zero bifurcation with $\mathbb{Z}_{2}$–symmetry [27, pp. 400, Case c=1c=-1]. The next result deals with the description of all bifurcation curves passing through the bifurcation point $(b,c)=(1,\pm 1)$ and that characterises a DZ bifurcation of $f_{(0,b,c)}$ with $\mathbb{Z}_{2}(\kappa)$–symmetry.
We focus the proof on the analysis around the bifurcation parameter $(b,c)=\left(1,1\right)$; the analysis close to $(b,c)=\left(1,-1\right)$ has a similar treatment (reversing time).
@@
Regarding to the vector field $f_{(0,b,c)},$ the equilibrium $E_{1}\equiv E_{2}\equiv E_{3}$ undergoes a DZ bifurcation with $\mathbb{Z}_{2}(\kappa)$–symmetry at $(b,c)=(1,1).$ The global representations of the transition/bifurcation curves in the parameters space $(b,c)\in\mathbb{R}\times\mathbb{R}^{+}$ are as follows (schematic lines for $a=0$ have been plotted in Figures 1 and 3):
@@
    The equilibrium $E_{1}$ undergoes a supercritical (with respect to $b$) Hopf bifurcation along part of the parabola:
@@
    The equilibria $E_{2}$ and $E_{3}$ undergo a subcritical (with respect to $b$) Hopf bifurcation along the algebraic curves:
@@
Remark 3.5.
For every pair $(b,c)\in\mathbb{R}\times\mathbb{R}^{+}$, the origin is always an equilibrium point. When $b<0$, two additional nontrivial equilibria appear; they move away from the origin and diverge to infinity as $b\to 0^{-}$. For $0<b<1$, the origin is the only equilibrium. Thus, although the number of equilibria changes as the line $b=0$ is crossed, this change results from equilibria escaping to (or emerging from) infinity rather than from a local bifurcation at the origin.
@@
Figure 1. Illustration of Theorem 3.4. Left: The bifurcation curves associated to (3.1) for Case A. Linear stability of $E_{1}$ for $f_{(0,b,c)}$. Right: Linear stability of $E_{2}$ and $E_{3}$ for $f_{(0,b,c)}$. The notation $T_{P}$, $T_{H}^{1}$, $T_{F}^{1}$, $T_{P}$ $T_{H}^{2,3}$, $T_{F}^{2,3}$ follows from Theorem 3.4 and sd, uf, un, sf, and sn refer to saddle, unstable focus, unstable node, stable focus and stable node.
@@
Theorem 3.4 generalizes the analysis of [8] and [21] by locating a codimension 2 bifurcation with $\mathbb{Z}_{2}(\kappa)$–symmetry at $(b,c)=(1,1)$ for the vector field $f_{(0,b,c)}.$ For the sake of completeness, we present local numerical phase portraits on Table 1 of Appendix. All of them agree well with the theory.
@@
    We prove the existence of a Pitchfork bifurcation for $b=1$ using the normal form truncated at order 3 [27, Section 20.1E]. Let $b=1+\mu$, where $\mu\in\mathbb{R}$ is an additional bifurcation parameter. For $\mu=0$, the Jacobian matrix (3.5) associated with $E_{1}$ admits eigenvalues: 0 and $\frac{c^{2}-1}{c}$, $c\neq 0,1$. Define the diffeomorphism $T:\mathbb{R}^{2}\to\mathbb{R}^{2}$ as
    $\mathbf{x}\equiv(x,y)\mapsto\left(\frac{c^{2}y-x}{c^{2}-1},-\frac{y-x}{c^{2}-1}\right).$
    In the new coordinates $\mathbf{x}=(x,y),$ where $\mathbf{0}=(0,0)$, let
@@
    We aim to find a center manifold in the form of
    $h(x,\mu):={\alpha_{0}}\mu^{2}+\alpha_{1}x^{2}+\alpha_{2}x\mu+\dots.$
    Since the center manifold must satisfy the equation (cf. [27, Equation 3.2.7])
@@
    meaning that there is a Pitchfork bifurcation at $\mu=0$. If $c>1$, then the bifurcation is supercritical with respect to $\mu$ and also with respect to $b$ in the original equation, as depicted in Figure 2. The analysis for the case $c>1$ is different from the case $0<c<1$, although the argument runs along the same lines.
@@
    Let $b=c^{2}$ and $|c|<1$. The Jacobian matrix of $f_{(0,b,c)}$ at $E_{1}$ has a pair of purely imaginary eigenvalues $i\sqrt{1-c^{2}}$ with 1<c<1.-1<c<1. Let $b=c^{2}+\mu$ where $\mu\in\mathbb{R}$ is a bifurcation parameter. The eigenvalues associated with the Jacobian matrix at $E_{1}$ for $\mu\neq 0$
@@
    $\begin{pmatrix}u\\ v\end{pmatrix}=\begin{pmatrix}\dfrac{1}{2}&\dfrac{c}{2\sqrt{1-c^{2}}}\\ 0&-\dfrac{1}{2c\sqrt{1-c^{2}}}\end{pmatrix}\begin{pmatrix}x\\ y\end{pmatrix}$,
@@
    The first Lyapunov coefficient depends on $c$ and is given by:
    $\ell(c)=\dfrac{1}{16\omega}\left(R_{1}+\omega R_{2}\right),$